# Discontinuous integro-differential equations and sliding mode control

### v2

@@ Abstract
The paper deals with analysis and design of sliding mode control systems modeled by finite-dimensional integro-differential equations. Filippov method and equivalent control approach are extended to a class of nonlinear discontinuous integro-differential equations. Sliding mode control algorithms are designed for distributed input delay systems and for systems modeled by differential equations in Banach spaces. The obtained results are illustrated by numerical examples.
@@ 1.1 State of the art
Sliding Mode Control (SMC) is a robust control methodology, which uses discontinuous feedback laws for rejection of matched disturbances. The design and analysis of SMC is well-developed for finite-dimensional systems [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book]. Mathematical models of finite-dimensional SMC systems are ill-posed discontinuous ODEs (Ordinary Differential Equations). Filippov method [Filippov1988:Book], [Cortes2008:CSM] is aimed at regularization of discontinuous ODEs (Ordinary Differential Equations). The method consists in a minimal (in a set-theoretic sense) extension of an ill-posed discontinuous ODE to a well-posed ODI (Ordinary Differential Inclusion). The Filippov regularization has a rather simple mathematical formulation [Filippov1988:Book] based on a notion of the set of Lebesgue measure zero. Filippov method is the main tool for regularization of a SMC system [Filippov1988:Book], [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book], [PolyakovFridman2014:JFI], [Utkin_etal2020:Book] while an analysis of the closed-loop dynamics of the SMC system is usually based on the so-called equivalent control method [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book].
Formally, the finite-dimensional Filippov method is inapplicable to discontinuous infinite-dimensional systems in abstract spaces [DaleckyKrein1974:Book] (e.g., due to the absence of a notion of the set of Lebesgue measure zero in an abstract space). Traditionally (see, e.g., [OrlovUtkin1985:DAN], [Orlov_elal2004:IJC], [Orlov2008:Book] and references therein), SMC design for infinite-dimensional systems is based on an equivalent control method. Well-posedness analysis of infinite-dimensional SMC systems requires rather complicated constructions [Levaggi2002:DIE], [Pisano_etal2011:SIAM_JCO], [Levaggi2013:DCDS], [Orlov2020:Book]. Infinite-dimensional differential inclusion obtained from discontinuous infinite-dimensional differential equation in an abstract space (like in [Mohet_etal2023:SCL]) is a non-trivial object for research [Tolstonogov2000:Book]. Some generalized solutions to discontinuous functional differential equations of retarded type are introduced on [KolmanovskiiMyshkis1992:Book, page 58]. The corresponding regularization procedure consists in a construction of a convex hull of limits of all possible convergent sequences in a function space. Applicability of such non-constructive procedure to analysis and design of SMC systems seems questionable too.
Integro-Differential Equations (IDEs) [Lakshmikantham_etal1995:Book], [OReganMeehan1998:Book] model various finite and infinite dimensional systems, such as epidemics [KermackMcKendrick1991:BMB], dispersive waves [Debnath2005:Book], systems with delays [Hale1977:Book], [Fridman2014:Book], [Orlov2020:Book], and others. Input-output dynamics of some systems modeled by PDEs (Partial Differential Equations) can also be described by IDEs (see Section 1.2). Being a model of (possibly) infinite-dimensional system, an IDE is usually defined by functions in finite-dimensional spaces (see, e.g., [Lakshmikantham_etal1995:Book], [OReganMeehan1998:Book]). To the best of author’s knowledge, finite-dimensional Filippov method and SMC theory has never been developed for IDEs.
@@ 1.2 Motivation example: SMC design for a heat system
Let us consider a control system modeled by the PDE
@@
A control system similar to (1) also appears in the case of the boundary control design using the Dirichlet lifting [PrieurTrelat2019:TAC], [KatzFridman2021:EJC].
@@
This is a SMC provided that $q$ is properly selected [Orlov_elal2004:IJC], [Orlov2020:Book].
The infinite-dimensional closed-loop control system (1) – (6) is ill-posed in the SMC case. A regularization of this discontinuous PDEs with SMC is a nontrivial problem (see, e.g., [Orlov2020:Book, pages 24-28], [Mohet_etal2023:SCL]). However, the considered infinite-dimensional control system has the scalar input $u$ and the scalar output $y$. The discontinuous feedback $u$ is a function of output variable only. For such infinite-dimensional system, the well-posedness analysis can be reduced to a study of a finite-dimensional IDE that represents the input-output dynamics:
@@
where $y(0)=\int^{1}_{0}\xi(z)x^{0}(z)dz$, $b=\int^{1}_{0}\xi(z)\beta(z)dz$, and $p,\Phi$ are some decreasing functions (see Section 4). If the control law $u$ is defined by the formula (6), then the input-output dynamics of the system (1) - (6) should satisfy the discontinuous IDE (7), (6). Due to discontinuity of $u$ on $y$, this IDE is ill-posed [Lakshmikantham_etal1995:Book], [OReganMeehan1998:Book]. However, the IDE is defined in terms of finite-dimensional functions, which admit Filippov regularization [Filippov1988:Book]. This paper extends the Filippov method to a class of finite-dimensional IDEs, which covers the IDE (7), (6) as a particular case. A regularization of the IDE (7), (6) allows the original discontinuous PDE (1) – (6) to be easily regularized (see, e.g., [Polyakov2025:CDC]). Such an IDE-based approach to SMC design and analysis is applicable to a rather large class of systems modeled by differential equations in Banach spaces
@@ 1.3 Contributions of the paper
  • A SMC design for some control systems with distributed input delay is presented.
  • An IDE-based method of SMC design and analysis for systems modeled by differential equations in Banach spaces is developed and illustrated on the system (1)-(6).
@@ 1.4 Organization of the paper
The paper is organized as follows. First, Filippov method for discontinuous ODEs is briefly surveyed. The surveyed results are utilized in the next section, where an extension of Filippov method to IDEs is proposed. After that, the developed theoretical results are applied for SMC design and analysis of finite-dimensional systems with distributed input delay and of infinite-dimensional systems modeled by differential equations in Banach spaces. Finally, concluding remarks are given.
@@ 2.1 Some classes of functions
For $1\!\leq\!p\!\leq\!\infty$, $L^{p},$ denotes the well-known Lebesgue space. A function $(t,x)\in\mathbb{R}\times\mathbb{R}^{n}\mapsto f(t,x)\in\mathbb{R}^{n}$ is said to satisfy
  • Carathéodory Condition if it is measurable in $t$ for any $x$, continuous in $x$ for almost any $t$ and for any $\alpha,r>0$ there exists $m\in L^{1}((-\alpha,\alpha);\mathbb{R})$ such that $\left\|f(t,x)\right\|\leq m(t)$ for almost all $t\!\in\!(-\alpha,\alpha)$, for all $x\!\in\!\mathcal{B}(r)$, where $\mathcal{B}(r)$ is an open ball of the radius $r>0$.
@@ 2.2 Well-posed ODEs
defines a rule of an evolution of the state $x(t)\in\mathbb{R}^{n}$ of the system in time $t\in\mathbb{R}$. Let $\mathcal{I}$ be one of the following intervals: $[a,b]$, $(a,b)$, $[a,b)$ or $(a,b]$, where $a,b\in[-\infty,+\infty]$.
@@
A function $x\in C^{1}(\mathcal{I};\mathbb{R}^{n})$ is said to be a classical solution of ODE (11) if it satisfies the equation (11) everywhere on $\mathcal{I}$.
@@
For the proofs of the classical theorem surveyed in this subsection we refer the reader, for example, to [CoddingtonLevinson1955:Book], [Filippov1988:Book].
@@
An absolutely continuous function $x:\mathcal{I}\subset\mathbb{R}^{n}\to\mathbb{R}^{n}$ is said to be a Carathéodory (or strong) solution of ODE (11) if it satisfies the equation (11) almost everywhere on the time interval $\mathcal{I}$.
@@
f $f$ satisfies Carathéodory Condition then the IVP (11), (12) has a Carathéodory solution defined, at least, locally in time.
@@
Theorem 3 (Continuity of solutions on initial data [CoddingtonLevinson1955:Book], page 58).
@@ 2.3 Regularization of ill-posed ODE by Filippov method
Systems with discontinuous in $x$ right-hand sides may have Carathéodory solutions in some cases [Liberzon2003:Book]. However, SMC theory deals with control systems modeled by ill-posed ODEs with discontinuous right-hand sides [Utkin_etal2009:Book, Liberzon2003:Book, GeligLeonovYakubovich2004:Book].
@@
To define a solution of the system (11) with discontinuous (on $x$) right-hand side, a special regularization procedure known as the Filippov method [Filippov1988:Book] can be utilized. It extends an ill-posed ODE to a well-posed ODI
@@
where $K[f]:\mathbb{R}\times\mathbb{R}^{n}\to 2^{\mathbb{R}^{n}}$ is defined as [Filippov1988:Book, page 85]:
@@
where $\overline{\mathrm{co}}$ denotes the convex closure, and the intersections are taken over all sets $M\subset\mathbb{R}^{n}$ of measure zero and all $\delta>0$. The set-valued mapping $(t,x)\to K[f](t,x)$ is nonempty-valued, compact-valued, convex-valued and upper semi-continuous in $x$ for almost all $t$ provided that $f$ is a measurable function (see, e.g., [Filippov1988:Book]). Filippov regularization (15) does not modify continuous in $x$ function $f$.
@@
Continuity of Filippov solutions on initial data [Filippov1988:Book, page 98] repeats the classical result if solutions of (14) are unique.
@@ 2.4 Uniqueness of Filippov solutions of ODEs
Lipschitz condition does not hold for discontinuous (on state variable) ODEs. Filippov solution of a discontinuous ODE is understood as a solution of an ODI. The uniqueness analysis for solutions of the ODI is a non-trivial problem in the general case [Filippov1988:Book], [Tolstonogov2000:Book]. However, One-Sided Lipschitz Condition is still applicable for uniqueness analysis of Filippov solutions.
Theorem 7 (Uniqueness of Filippov solutions, [Filippov1988:Book], §10).
@@
where the set-valued function $\mathcal{N}:\mathcal{S}\to 2^{\{1,2,...,N\}}$ indicates the domains, which have a common boundary point $x\in\mathcal{S}$, i.e., $\mathcal{N}(x)=\left\{j\in\{1,2,...,N\}:x\in\partial G^{j}\right\}.$
Let $\mathcal{C}_{\Omega}(x)\subset\mathbb{R}^{n}$ denote the (Bouligand) tangent cone to the set $\Omega\subset\mathbb{R}^{n}$ at $x\in\mathbb{R}^{n}$ defined as (see., e.g. [Nagumno1942:PPMSJ])
@@
Tangent cones characterize positively invariant sets [^2] A set $M\!\subset\!\mathbb{R}^{n}$ is positively invariant if $x(t^{0})\!\in\!M\Rightarrow x(t)\!\in\!M,t\!\geq\!t_{0}.$ of dynamical systems (see, e.g., [Clarke_etal1995:JDCS], [Blanchini1999:Aut], [AubinEkeland1984:Book] and references therein).
@@
Let us denote $F_{\mathcal{S}}(t,x)=K[f](t,x)\cap\mathcal{C}_{\mathcal{S}}(x),t\in\mathbb{R},x\in\mathcal{S}$. The following result is partially inspired by [Filippov1988:Book, §10].
@@
If the set $F_{\mathcal{S}}$ is a singleton $F_{\mathcal{S}}(t,x)=\{f^{0}(t,x)\}$ for all $x\in\mathcal{S}$, for almost all $t\in\mathbb{R}$ and for some locally one-sided Lipschitz continuous function $f^{0}$, then, under Assumption 1, the IVP (11), (12) has a Filippov solution being unique, at least, locally in the forward time.
@@
The system (16) satisfies the above theorem with $F_{\mathcal{S}}=\{0\}$. The relay SMC system is the conventional example of discontinuous ODE satisfying the above theorem [Utkin1992:Book], [Filippov1988:Book].
@@
The simplest example of discontinuous ODE with unique solutions can be discovered in [Utkin1992:Book] for the system
@@
Let $f$ be piece-wise one-sided Lipschitz continuous and $F_{\mathcal{S}}(t,x)$$\emptyset$ for all $x\in\mathcal{S}$, for almost all $t\in\mathbb{R}$.
@@
  • $F_{\mathcal{S}}(t,x)$$\emptyset$, $\forall x\!\in\!\mathcal{S}\backslash\mathcal{M}$, for almost all $t\!\in\!\mathbb{R}$;
@@
Assumption 3 ensures that $\mathcal{M}$ is the sliding set and the (switching) set $\mathcal{S}\backslash\mathcal{M}$ is crossed by trajectories without sliding.

### v3

@@ Abstract
The paper deals with analysis and design of sliding mode control systems modeled by finite-dimensional integro-differential equations. Filippov method and equivalent control approach are extended to a class of nonlinear discontinuous integro-differential equations and to a class of control systems modeled by infinite-dimensional differential equations in Banach spaces. Sliding mode control algorithms are designed for distributed input delay systems and for a heat control system.
@@ 1.1 State of the art
Sliding Mode Control (SMC) is a robust control methodology, which uses discontinuous feedback laws for rejection of matched disturbances. The design and analysis of SMC is well-developed for finite-dimensional systems [49], [14], [46]. Mathematical models of finite-dimensional SMC systems are ill-posed discontinuous ODEs (Ordinary Differential Equations). Filippov method [17], [11] is aimed at regularization of discontinuous ODEs. The method consists in a minimal (in a set-theoretic sense) extension of an ill-posed discontinuous ODE to a well-posed ODI (Ordinary Differential Inclusion). The Filippov regularization has a rather simple mathematical formulation [17] based on the notion of a set of Lebesgue measure zero. Filippov method is the main tool for regularization of a SMC system [17], [49], [14], [46], [43], [50] while an analysis of the closed-loop dynamics of the SMC system is usually based on the so-called equivalent control method [49], [14], [46].
Formally, the finite-dimensional Filippov method is inapplicable to discontinuous infinite-dimensional systems in abstract spaces [12] (e.g., due to the absence of a notion of the set of Lebesgue measure zero in an abstract space). Traditionally (see, e.g., [36], [35], [37] and references therein), SMC design for infinite-dimensional systems is based on an equivalent control method. Well-posedness analysis of infinite-dimensional SMC systems requires rather complicated constructions [27], [41], [28], [38]. Infinite-dimensional differential inclusion obtained from discontinuous infinite-dimensional differential equation in an abstract space (like in [31]) is a non-trivial object for research [47]. Some generalized solutions to discontinuous functional differential equations of retarded type are introduced on [24, page 58]. The corresponding regularization procedure consists in a construction of a convex hull of limits of all possible convergent sequences in a function space. Applicability of such non-constructive procedure to analysis and design of SMC systems seems questionable too.
Integro-Differential Equations (IDEs) [26], [34] model various finite and infinite dimensional systems, such as epidemics [23], dispersive waves [13], systems with delays [21], [18], [38], and others. Input-output dynamics of some systems modeled by PDEs (Partial Differential Equations) can also be described by IDEs (see Section 1.2). Being a model of (possibly) infinite-dimensional system, an IDE is usually defined by functions in finite-dimensional spaces (see, e.g., [26], [34]). To the best of author’s knowledge, finite-dimensional Filippov method and SMC theory has never been developed for IDEs.
@@ 1.2 Motivation example: SMC design for a heat system
Let us consider a control system modeled by the heat PDE
@@
A control system similar to (1) also appears in the case of the boundary control design using the Dirichlet lifting [45], [22].
@@
This is a SMC provided that $q$ is properly selected [35], [38].
The infinite-dimensional closed-loop control system (1) – (6) is ill-posed in the SMC case. A regularization of this discontinuous PDEs with SMC is a nontrivial problem (see, e.g., [38, pages 24-28], [31]). However, the considered infinite-dimensional control system has the scalar input $u$ and the scalar output $y$. The discontinuous feedback $u$ is a function of output variable only. For such infinite-dimensional system, the well-posedness analysis can be reduced to a study of a finite-dimensional IDE that represents the input-output dynamics:
@@
where $y(0)=\int^{1}_{0}\xi(z)x^{0}(z)dz$, $b=\int^{1}_{0}\xi(z)\beta(z)dz$, and $p,\Phi$ are some continuous functions (see Example 8 in Section 4).
If the control law $u$ is defined by the formula (6), then the input-output dynamics of the system (1) - (6) should satisfy the discontinuous IDE (7), (6). Due to discontinuity of $u$ on $y$, this IDE is ill-posed [26], [34]. However, the IDE is defined in terms of finite-dimensional functions, which admit Filippov regularization [17]. This paper extends the Filippov method to a class of finite-dimensional IDEs, which covers the IDE (7), (6) as a particular case. A regularization of the IDE (7), (6) allows the original discontinuous PDE (1) – (6) to be easily regularized (see, e.g., [44]). Such an IDE-based approach to SMC design and analysis is applicable to a rather large class of systems modeled by differential equations in Banach spaces.
@@ 1.3 Contributions of the paper
  • A SMC design for control systems modeled by linear IDEs is presented.
  • An IDE-based SMC design and analysis for some control systems modeled by differential equations in Banach spaces is developed and illustrated on the system (1)-(6).
@@ 1.4 Organization of the paper
The paper is organized as follows. First, Filippov method for discontinuous ODEs is briefly surveyed. The surveyed results are utilized in the next section, where an extension of Filippov method to IDEs is proposed. After that, the obtained theoretical results are applied to SMC design and analysis for finite-dimensional systems with distributed input delay and for infinite-dimensional systems modeled by differential equations in Banach spaces. Finally, concluding remarks are given.
@@ 2.1 Some classes of functions
For $1\!\leq\!p\!\leq\!\infty$, $L^{p}$ denotes the well-known Lebesgue space. A function $(t,x)\in\mathbb{R}\times\mathbb{R}^{n}\mapsto f(t,x)\in\mathbb{R}^{n}$ is said to satisfy
  • Carathéodory Condition if $f$ is measurable in $t$ for all $x$, continuous in $x$ for almost all $t$ and for any $\alpha,r>0$ there exists $m\in L^{1}((-\alpha,\alpha);\mathbb{R})$ such that $\left\|f(t,x)\right\|\leq m(t)$ for almost all $t\!\in\!(-\alpha,\alpha)$, for all $x\!\in\!\mathcal{B}(r)$, where $\mathcal{B}(r)$ is an open ball of the radius $r>0$, $\|\cdot\|$ - Euclidean norm.
@@ 2.2 Well-posed ODEs
defines a rule of an evolution of the state $x(t)\in\mathbb{R}^{n}$ of the system in time $t\in\mathbb{R}$. Let $\mathcal{I}$ be one of the following intervals: $[a,b]$, $(a,b)$, $[a,b)$ or $(a,b]$, where $a,b\in[-\infty,+\infty],a<b$.
@@
A function $x\in C^{1}(\mathcal{I};\mathbb{R}^{n})$ is said to be a classical solution of ODE ($1$) if it satisfies the equation (11) everywhere on $\mathcal{I}$.
@@
For the proofs of the classical theorems surveyed in this subsection we refer the reader, for example, to [10], [17].
@@
An absolutely continuous function $x:\mathcal{I}\subset\mathbb{R}^{n}\to\mathbb{R}^{n}$ is said to be a Carathéodory (or strong) solution of ODE ($1$) if it satisfies the equation (11) almost everywhere on the time interval $\mathcal{I}$.
@@
If $f$ satisfies Carathéodory Condition then the IVP (11), (12) has a Carathéodory solution defined, at least, locally in time.
@@
Theorem 3 (Continuity of solutions on initial data [10], page 58).
@@ 2.3 Regularization of ill-posed ODE by Filippov method
Systems with discontinuous in $x$ right-hand sides may have Carathéodory solutions in some cases [30]. However, SMC theory deals with control systems modeled by ill-posed ODEs with discontinuous right-hand sides [48, 30, 19].
@@
To define a solution of the system (11) with discontinuous (on $x$) right-hand side, a special regularization procedure known as the Filippov method [17] can be utilized. It extends an ill-posed ODE to a well-posed ODI
@@
where $K[f]:\mathbb{R}\times\mathbb{R}^{n}\to 2^{\mathbb{R}^{n}}$ is defined as [17, page 85]:
@@
where $\overline{\mathrm{co}}$ denotes the convex closure, and the intersections are taken over all sets $M\subset\mathbb{R}^{n}$ of measure zero and all $\delta>0$ and $2^{\mathbb{R}^{n}}$ denotes the power set of $\mathbb{R}^{n}$. The set-valued mapping $(t,x)\to K[f](t,x)$ is nonempty-valued, compact-valued, convex-valued and upper semi-continuous in $x$ for almost all $t$ provided that $f$ is a measurable function (see, e.g., [17]). Notice that Filippov regularization (15) does not modify continuous in $x$ function $f$.
@@
Continuity of Filippov solutions on initial data [17, page 98] repeats the classical result if solutions of (14) are unique.
@@ 2.4 Uniqueness of Filippov solutions of ODEs
Lipschitz condition does not hold for discontinuous (on state variable) ODEs. Filippov solution of a discontinuous ODE is understood as a solution of an ODI. The uniqueness analysis for solutions of the ODI is a non-trivial problem in the general case [17], [47]. However, One-Sided Lipschitz Condition is still applicable for uniqueness analysis of Filippov solutions.
Theorem 7 (Uniqueness of Filippov solutions, [17], §10).
@@
where, $\mathcal{S}=\bigcup_{i=1}^{N}\partial G^{j}$ is the discontinuity set of $f$, but the set-valued function $\mathcal{N}:\mathcal{S}\to 2^{\{1,2,...,N\}}$ indicates the continuity domains $G^{j}$, which have a common boundary point $x\in\mathcal{S}$, i.e., $\mathcal{N}(x)=\left\{j\in\{1,2,...,N\}:x\in\partial G^{j}\right\}.$
Let $\mathcal{C}_{\Omega}(x)\subset\mathbb{R}^{n}$ denote the (Bouligand) tangent cone to the set $\Omega\subset\mathbb{R}^{n}$ at $x\in\mathbb{R}^{n}$ defined as (see., e.g. [32])
@@
Tangent cones characterize positively invariant sets [^2] A set $M\!\subset\!\mathbb{R}^{n}$ is positively invariant if $x(t^{0})\!\in\!M\Rightarrow x(t)\!\in\!M,t\!\geq\!t_{0}.$ of dynamical systems (see, e.g., [9], [7], [2] and references therein).
@@
The following result is partially inspired by [17, §10].
@@
Let
$F_{\mathcal{S}}(t,x)=K[f](t,x)\cap\mathcal{C}_{\mathcal{S}}(x),\quad t\in\mathbb{R},\quad x\in\mathcal{S}.\vskip-2.84526pt$ (19)
If the set $F_{\mathcal{S}}$ is a singleton, i.e., $F_{\mathcal{S}}(t,x)=\{f^{0}(t,x)\}$ for all $x\in\mathcal{S}$, for almost all $t\in\mathbb{R}$, with for some locally one-sided Lipschitz continuous function $f^{0}$, then, under Assumption 1, the IVP (11), (12) has a Filippov solution being unique, at least, locally in the forward time.
@@
The system (16) satisfies the above theorem with $F_{\mathcal{S}}=\{0\}$. The relay SMC system is the conventional example of discontinuous ODE satisfying the above theorem [49], [17].
@@
A simple example of discontinuous ODE with unique solutions can be discovered in [49] for the SMC system
@@
Let $f$ be piece-wise one-sided Lipschitz continuous and $F_{\mathcal{S}}(t,x)=\emptyset$ for all $x\in\mathcal{S}$, for almost all $t\in\mathbb{R}$.
@@
  • $F_{\mathcal{S}}(t,x)=\emptyset$, $\forall x\!\in\!\mathcal{S}\backslash\mathcal{M}$, for almost all $t\!\in\!\mathbb{R}$;
@@
Assumption 3 ensures that $\mathcal{M}$ is the sliding set and the switching set $\mathcal{S}\backslash\mathcal{M}$ is crossed by trajectories without sliding.